JustSaiyan:

Solve the compound inequality. Show your steps. 3x + 2 ≤ 10 OR 3x + 2 ≥ 15

MARC:

Subtract both sides by 2 Subtract both sides by 2 \(3x+2-2\le10-2\) or \(3x+2-2\ge15-2\) \(3x\le8\) \(3x\ge13\) Divide both sides by 3 Divide both sides by 3 \(\frac{3}{3}x\le\frac{8}{3}\) \(\frac{3}{3}\ge\frac{13}{3}\)

MARC:

\(x\le?\) or \(x\ge?\) solve it @JustSaiyan